A bond issued by the national government pays 1000 Php at the end of each year for 6 years, plus an additional 10,000 Php when the bond matures at the end of 6 years. What is the maximum payment for this bond if the existing opportunity cost of funds is 10%?


1
Expert's answer
2021-08-24T14:09:06-0400

Solution:

The maximum payment = "PMT\times [1 - \frac{1}{(1 +r)^{n} }\div r]" PMT x [1 – (1/1+r)n)/r]


"=1,000\times [1 - \frac{1}{(1 +0.1)^{6} }\div 0.1]"


"=1,000\times0.57954 = 579.54"


The maximum payment for this bond is = 579.54 Php